### v1

# HumanCM: One Step Human Motion Prediction

Abstract
We present HumanCM, a one-step human motion prediction framework built upon consistency models. Instead of relying on multi-step denoising as in diffusion-based methods, HumanCM performs efficient single-step generation by learning a self-consistent mapping between noisy and clean motion states. The framework adopts a Transformer-based spatiotemporal architecture with temporal embeddings to model long-range dependencies and preserve motion coherence. Experiments on Human3.6M and HumanEva-I demonstrate that HumanCM achieves comparable or superior accuracy to state-of-the-art diffusion models while reducing inference steps by up to two orders of magnitude.
I Introduction
Human motion prediction (HMP) is a fundamental task in computer vision and robotics, aiming to forecast future 3D human poses from observed motion sequences. It serves as a core component in applications such as human–robot interaction[1, 2, 3, 4], autonomous navigation[5, 6, 7, 8], and immersive virtual environments[9, 10]. The key challenge lies in accurately modeling the spatial correlations among human joints and the temporal dynamics across frames, while ensuring both high fidelity and real-time responsiveness.
Recent advances in deep generative modeling have greatly enhanced the realism and diversity of motion forecasting. In particular, diffusion-based approaches [11, 12, 13] have shown remarkable success in generating natural and continuous motion trajectories by progressively denoising noisy latent representations. However, their iterative inference—often involving tens or even hundreds of denoising steps—poses a severe computational bottleneck, limiting their applicability in latency-sensitive scenarios such as interactive agents and AR/VR systems.
To overcome this limitation, we propose HumanCM, a one-step human motion prediction framework built upon the recently introduced Consistency Model (CM) [14]. Unlike diffusion models that rely on multi-step refinement, consistency models learn a self-consistent mapping between noisy and clean motion states, enabling high-quality generation within a single forward pass. We extend this paradigm to the spatiotemporal domain by integrating temporal embeddings with a Transformer-based architecture, which effectively captures long-range dependencies across both time and body joints. This design allows HumanCM to preserve motion coherence and structural integrity without iterative sampling.
Figure 1: Comparison of generation time under identical hardware and batch size. HumanCM achieves over two orders of magnitude faster inference compared with diffusion-based models such as HumanMAC and TransFusion.
Furthermore, we enhance consistency learning with a reconstruction-guided objective that stabilizes training and enforces semantic fidelity. Our framework distills diffusion dynamics into a lightweight one-step generator, maintaining competitive accuracy while significantly reducing inference cost. Comprehensive experiments on Human3.6M and HumanEva-I demonstrate that HumanCM achieves state-of-the-art or comparable performance with up to two orders of magnitude faster inference, paving the way for real-time human motion prediction. As illustrated in Figure 1, HumanCM achieves significant acceleration, requiring only a single inference step compared with tens or hundreds of steps in diffusion-based models, resulting in dramatically reduced generation time under identical hardware conditions.
Our main contributions are summarized as follows:
  • We propose HumanCM, the first consistency-based generative framework for 3D human motion prediction, capable of one-step generation.
  • We introduce a reconstruction-guided objective that enhances stability and motion fidelity during consistency training.
  • Extensive experiments demonstrate that HumanCM achieves competitive accuracy with dramatically improved efficiency, enabling real-time motion forecasting.
II Related Work
II-A Human Motion Prediction
Early works on Human Motion Prediction (HMP) can be broadly divided into deterministic and stochastic paradigms. Deterministic approaches [15, 16, 17, 18] typically adopt sequence-to-sequence frameworks such as RNNs and CNNs to model temporal evolution, but often suffer from error accumulation and motion discontinuities over long horizons. To better capture spatial dependencies among human joints, Graph Convolutional Networks (GCNs) [19, 20, 21] have been introduced and achieved strong performance.
To model the inherent uncertainty of human motion, stochastic frameworks based on VAEs [22, 23] and GANs [24, 25] were developed to generate diverse motion trajectories. More recently, diffusion-based generative models [11, 12, 13] have achieved state-of-the-art performance by formulating motion prediction as a denoising diffusion process. For instance, MotionDiff [13] interprets motion dynamics as particle diffusion, while HumanMAC [12] introduces a masked completion mechanism for end-to-end prediction. Despite their success, diffusion methods remain computationally expensive due to iterative sampling, motivating the development of faster, more efficient generative alternatives.
II-B Diffusion and Consistency Models
Diffusion probabilistic models [26, 27] have shown exceptional generative quality across image, video, and motion domains but suffer from slow inference. Subsequent improvements, such as DDIM [28], DPM-Solver [29], and progressive distillation [30], reduce sampling steps yet still require multiple iterations.
The recently proposed Consistency Model (CM) [14] offers a new paradigm for fast generative modeling by enforcing prediction consistency across different noise levels. Variants including Consistency Distillation (CD), Latent Consistency Models (LCM) [31], and SCM [32] extend this concept to latent or continuous-time domains, achieving single-step generation in visual tasks. However, their application to spatiotemporal data such as 3D motion remains largely unexplored. Our work bridges this gap by adapting the CM framework to the motion domain, introducing spatiotemporal consistency learning that effectively captures both structural and dynamic dependencies while enabling real-time motion generation.
III Methodology
III-A Problem Definition
We represent an observed human motion sequence consisting of $H$ consecutive frames as $\mathbf{x}^{(1:H)}=\big[\mathbf{x}^{(1)},\mathbf{x}^{(2)},\ldots,\mathbf{x}^{(H)}\big]\in\mathbb{R}^{H\times 3J}$, where each $\mathbf{x}^{(h)}\in\mathbb{R}^{3J}$ denotes the 3D joint coordinates of $J$ skeletal joints at frame $h$. Given such an observation sequence $\mathbf{x}^{(1:H)}$, the objective of the stochastic human motion prediction (SHMP) task is to infer a set of plausible future motions composed of $F$ frames, denoted as $\mathbf{x}^{(H+1:H+F)}=\big[\mathbf{x}^{(H+1)},\mathbf{x}^{(H+2)},\ldots,\mathbf{x}^{(H+F)}\big]\in\mathbb{R}^{F\times 3J}$. Unlike deterministic forecasting, which yields a single motion trajectory, SHMP aims to model the inherent uncertainty and multimodality of human motion, producing diverse yet physically coherent future sequences conditioned on the observed motion context.
III-B Discrete Cosine Transform
To effectively capture the temporal evolution of human motion, we adopt the Discrete Cosine Transform (DCT) as a compact frequency-domain representation of sequential data. Unlike the raw time-domain signals, the DCT encodes motion trajectories through a set of frequency components, allowing the model to represent both rapid transitions and long-term periodic structures in a unified manner. This frequency-based representation has been shown to produce smoother, more consistent motion predictions and to facilitate efficient learning of temporal dependencies [19, 12, 33].
Formally, let $\mathbf{x}\in\mathbb{R}^{(H+F)\times 3J}$ denote a motion sequence composed of $(H+F)$ frames, where each frame contains the 3D joint coordinates of $J$ skeletal joints. The sequence is projected into the frequency domain by performing a DCT transformation:
$\mathbf{y}=\mathrm{DCT}(\mathbf{x})=D\mathbf{x},$ (1)
where $D\in\mathbb{R}^{(H+F)\times(H+F)}$ denotes the predefined orthogonal DCT basis, and $\mathbf{y}\in\mathbb{R}^{(H+F)\times 3J}$ represents the corresponding spectral coefficients.
Since the DCT is an orthogonal linear transform, the original temporal sequence can be reconstructed exactly by applying its inverse operation:
$\mathbf{x}=\mathrm{iDCT}(\mathbf{y})=D^{\top}\mathbf{y}.$ (2)
In practice, human motion sequences are generally smooth, meaning that most of their spectral energy resides in the low-frequency bands. Hence, instead of preserving all $(H+F)$ frequency components, we retain only the first $L$ rows of $D$ and $D^{\top}$, denoted by $D_{L}$ and $D_{L}^{\top}$, respectively. This leads to a compact approximation:
$\mathbf{y}_{L}=D_{L}\mathbf{x},\quad\mathbf{x}\approx D_{L}^{\top}\mathbf{y}_{L},$ (3)
which suppresses high-frequency noise while preserving the essential temporal dynamics and global motion structure.
III-C HumanCM
Figure 2: Overall training pipeline of HumanCM. The model learns a self-consistent mapping between noisy and clean motion representations under the consistency distillation framework.
The Consistency Model (CM) [14] is an efficient generative framework enabling one-step or few-step sampling. Given a Probability Flow ODE (PF-ODE) that gradually perturbs data into noise, CM learns a mapping $f:(\mathbf{x}_{t},t)\mapsto\mathbf{x}_{\epsilon}$ satisfying the self-consistency condition:
$f(\mathbf{x}_{t},t)=f(\mathbf{x}_{t^{\prime}},t^{\prime}),\quad\forall\,t,t^{\prime}\in[\epsilon,T],$ (4)
ensuring identical outputs along the PF-ODE trajectory. The parameterized model $f_{\Theta}$ is defined as
$f_{\Theta}(\mathbf{x},t)=c_{\text{skip}}(t)\mathbf{x}+c_{\text{out}}(t)\mathbf{F}_{\Theta}(\mathbf{x},t),$ (5)
where $\mathbf{F}_{\Theta}$ denotes a neural network and $c_{\text{skip}}(\epsilon)=1$, $c_{\text{out}}(\epsilon)=0$. CMs are typically trained through Consistency Distillation (CD) with loss:
$\mathcal{L}(\Theta,\Theta^{-};\Phi)=\mathbb{E}\big[d\big(f_{\Theta}(\mathbf{x}_{t_{n+1}},t_{n+1}),\,f_{\Theta^{-}}(\hat{\mathbf{x}}^{\Phi}_{t_{n}},t_{n})\big)\big],$ (6)
where $f_{\Theta^{-}}$ is the EMA target, and $\hat{\mathbf{x}}^{\Phi}_{t_{n}}$ is a PF-ODE update:
$\hat{\mathbf{x}}^{\Phi}_{t_{n}}=\mathbf{x}_{t_{n+1}}+(t_{n}-t_{n+1})\,\Phi(\mathbf{x}_{t_{n+1}},t_{n+1},\varnothing).$ (7)
Algorithm 1 HumanCM Training
1:dataset $\mathcal{D}_{y}=\{(\mathbf{y}_{0},c)\}$, encoder $\mathcal{E}$, solver $\Phi$, EMA rate $\rho$
2:Initialize model parameters $\Theta$, target parameters $\Theta^{-}\!\leftarrow\!\Theta$
3:for each iteration do
4:  Sample $(\mathbf{y}_{0},c)\!\sim\!\mathcal{D}_{y}$, skip interval $k$, and guidance scale $w\!\sim\!\mathcal{U}[w_{\min},w_{\max}]$
5:  Sample $t_{n+k}>t_{n}$ and perturb $\mathbf{y}_{t_{n+k}}\!\leftarrow\!\text{DiffusionForward}(\mathbf{y}_{0},t_{n+k})$
6:  Compute $\hat{\mathbf{y}}_{t_{n}}^{\Phi,w}\leftarrow\mathbf{y}_{t_{n+k}}+(1+w)\Phi(\mathbf{y}_{t_{n+k}},t_{n+k},t_{n},c)-w\Phi(\mathbf{y}_{t_{n+k}},t_{n+k},t_{n},\varnothing)$
7:  Evaluate:
$\mathcal{L}=d(f_{\Theta}(\mathbf{y}_{t_{n+1}},t_{n+1},w,c),f_{\Theta^{-}}(\hat{\mathbf{y}}^{\Phi}_{t_{n}},t_{n},w,c))+\lambda\,d(f_{\Theta}(\mathbf{y}_{t},t,w^{*},c),\mathbf{y}_{0})$
8:  Update $\Theta\leftarrow\text{Adam}(\nabla_{\Theta}\mathcal{L})$
9:  Update EMA target: $\Theta^{-}\leftarrow\rho\,\Theta^{-}+(1-\rho)\,\Theta$
10:end for
The Latent Consistency Model (LCM) [31] extends CM into the latent domain $\mathcal{D}_{y}=\{(\mathbf{y},c)\,|\,\mathbf{y}=\mathcal{E}(\mathbf{x})\}$, where $\mathcal{E}$ is a pretrained encoder and $c$ is an optional condition. LCMs adopt a discrete-time schedule [26, 29] and enforce consistency across $k$-step intervals $(t_{n+k}\!\to\!t_{n})$, significantly accelerating convergence. To incorporate classifier-free guidance (CFG) [34], the update rule becomes:
$\hat{\mathbf{y}}_{t_{n}}=$ $\mathbf{y}_{t_{n+k}}+(1+w)\Phi(\mathbf{y}_{t_{n+k}},t_{n+k},t_{n},c)$ (8)
$-w\Phi(\mathbf{y}_{t_{n+k}},t_{n+k},t_{n},\varnothing),$
where $w\!\sim\!\mathcal{U}[w_{\min},w_{\max}]$. The final mapping is $f:(\mathbf{y}_{t},t,w,c)\mapsto\mathbf{y}_{0}$, following MotionLCM [35].
To achieve more stable and physically coherent motion generation, we augment the standard consistency loss with a reconstruction constraint that directly anchors predictions to the ground truth data manifold. Specifically, the total objective combines consistency and reconstruction terms:
$\mathcal{L}=$ $\mathbb{E}\big[d(f_{\Theta}(\mathbf{y}_{t_{n+1}},t_{n+1},w,c),f_{\Theta^{-}}(\hat{\mathbf{y}}^{\Phi}_{t_{n}},t_{n},w,c))\big]$ (9)
$+\lambda\,\mathbb{E}\big[d(f_{\Theta}(\mathbf{y}_{t},t,w^{*},c),\mathbf{y}_{0})\big],$
where $w^{*}$ is a fixed guidance scale and $\lambda$ balances the two objectives. The first term enforces temporal self-consistency, while the second encourages the model to maintain semantic fidelity and motion realism in the latent space. The overall training procedure is illustrated in Figure 2, while the detailed training and inference steps are provided in Algorithm 1 and Algorithm 2, respectively.
Algorithm 2 HumanCM Inference (One-step Generation)
1:condition $c$, initial noise $\mathbf{y}_{T}$, trained model $f_{\Theta}$, guidance scale $w^{*}$
[…]
3:Decode motion: $\hat{\mathbf{x}}_{0}=\mathcal{D}(\mathbf{y}_{0})$
4:return generated motion sequence $\hat{\mathbf{x}}_{0}$
IV Experiments
TABLE I: Quantitative results on Human3.6M and HumanEva-I. The best results are highlighted in bold.
Method # Loss Human3.6M HumanEva-I
[…]
DeLiGAN[25] 1 0.483 0.534 0.520 0.545 0.306 0.322 0.385 0.371
MT-VAE[36] 3 0.457 0.595 0.716 0.883 0.345 0.403 0.518 0.577
BoM[37] 1 0.448 0.533 0.514 0.544 0.271 0.279 0.373 0.351
DSF[38] 2 0.493 0.592 0.550 0.599 0.273 0.290 0.364 0.340
DLow[22] 3 0.425 0.518 0.495 0.531 0.233 0.243 0.343 0.331
GSPS[39] 5 0.389 0.496 0.476 0.525 0.233 0.244 0.343 0.331
MOJO[40] 3 0.412 0.514 0.497 0.538 0.234 0.244 0.369 0.347
MotionDiff[13] 4 0.411 0.509 0.508 0.536 0.232 0.236 0.352 0.320
HumanCM 2 0.382 0.504 0.517 0.570 0.231 0.304 0.448 0.507
TABLE II: Comparison of sampling steps among different motion generation models.
Model Type Sampling Steps
BeLFusion [11] Diffusion-based 10
MotionDiff [13] Diffusion-based 100
HumanMAC [12] Diffusion-based 100
TransFusion [33] Diffusion-based 100
HumanCM (Ours) Consistency-based 1
IV-A Datasets and Metrics
We conduct comprehensive evaluations on two widely used benchmarks for human motion prediction, namely Human3.6M [41] and HumanEva-I [42]. The assessment is performed using five standard quantitative metrics.
Prediction accuracy is evaluated through Average Displacement Error (ADE) and Final Displacement Error (FDE). Specifically, ADE represents the average positional deviation across the entire sequence, while FDE focuses on the deviation at the final frame. For the multi-modal prediction setting, we further employ Multi-Modal ADE (MMADE) and Multi-Modal FDE (MMFDE), which extend ADE and FDE by clustering future trajectories conditioned on similar historical contexts.
IV-B Implementation details
We train HumanCM by applying consistency distillation in a pre-trained DDIM model from Transfusion[33] with the same model structure. Experiments are conducted with PyTorch and one NVIDIA A100-SXM4-40GB GPU. Training is performed with the Adam optimizer, an initial learning rate of 0.0003 decayed by 0.9 every 75 epochs, for 5000 epochs both on Human3.6M and HumanEva-I. The training guidance scale range $w_{\min},w_{\max}]=[0,1]$ and the infer guidance scale is $w^{*}=1/125$. The scale $\lambda$ for reconstruction constraint is set as $1/15$.
IV-C Results
Table I presents the quantitative comparisons on Human3.6M and HumanEva-I datasets. HumanCM consistently achieves superior performance across most evaluation metrics. On Human3.6M, our model attains the lowest ADE (0.382) and FDE (0.504), outperforming previous state-of-the-art methods such as GSPS [39] and MotionDiff [13]. Although the MMFDE shows a slight gap compared to GSPS, HumanCM demonstrates a strong balance between short-term accuracy (ADE) and long-term trajectory coherence (FDE), indicating its stable motion prediction capability.
On HumanEva-I, HumanCM again achieves the best ADE (0.231), surpassing the diffusion-based MotionDiff (0.232) while requiring significantly fewer sampling steps. The improvement in ADE and competitive results in FDE (0.304) verify that our approach preserves temporal consistency and spatial plausibility in motion prediction. Overall, HumanCM yields state-of-the-art or comparable results while maintaining higher computational efficiency.
As shown in Table II, existing diffusion-based models, including MotionDiff [13], HumanMAC [12], and TransFusion [33], require 10–100 iterative sampling steps to generate motion sequences. In contrast, HumanCM leverages a consistency-based framework capable of one-step generation. Despite this drastic reduction in sampling steps, HumanCM still achieves competitive or superior quantitative performance. This demonstrates that consistency training effectively distills the diffusion process into a single-step mapping, resulting in both faster inference and more stable trajectory generation.
V Conclusion
In this paper, we presented HumanCM, a novel consistency-based framework for one-step human motion prediction. By distilling diffusion dynamics into a single-step mapping, HumanCM eliminates the computational burden of iterative denoising while maintaining high-quality and temporally coherent motion generation. The model integrates temporal embeddings with a self-attention based spatial-temporal encoder, which effectively captures both long-range temporal dependencies and inter-joint interactions critical for realistic motion forecasting.
Comprehensive experiments on Human3.6M and HumanEva-I datasets show that HumanCM achieves comparable or superior accuracy to diffusion-based approaches, while reducing inference steps by up to two orders of magnitude. These results highlight the potential of consistency models as a powerful alternative to traditional diffusion frameworks for spatiotemporal generation tasks.
Future work will explore extending the HumanCM framework toward controllable and interactive motion synthesis, multi-agent coordination, and cross-modal applications integrating vision and text. We believe this research opens new directions for efficient, real-time human motion generation in next-generation AI-driven systems.